Find particular integrals of the following partial differential equation to represent

surfaces passing through the following curve:

(𝑦 – 𝑧)𝑝 + (𝑧 – π‘₯)π‘ž = π‘₯ – 𝑦 ; 𝑧 = 0, 𝑦 = 2π‘₯


1
Expert's answer
2021-01-14T20:00:21-0500

"\displaystyle
(y - z)p + (z - x)q = x - y\\
\frac{\mathrm{d}x}{y - z} = \frac{\mathrm{d}y}{z - x} = \frac{\mathrm{d}x}{x - y}\\


\mathrm{d}x + \mathrm{d}y + \mathrm{d}z = 0\\


\int\mathrm{d}x + \int\mathrm{d}y + \int\mathrm{d}z = 0\\


x + y + z = c_1\\



\frac{\mathrm{d}y}{c_1 -2x - y} = \frac{\mathrm{d}x}{x - y}\\


(x - y)\mathrm{d}y + (2x + y - c_1)\mathrm{d}x = 0\\


x\mathrm{d}y + y\mathrm{d}x - y\mathrm{d}y + 2x\mathrm{d}x - c_1\mathrm{d}x = 0\\


\mathrm{d}(xy) - y\mathrm{d}y + 2x\mathrm{d}x - c_1\mathrm{d}x = 0\\


\int\mathrm{d}(xy) - \int y\mathrm{d}y + \int 2x\mathrm{d}x - \int c_1 \mathrm{d}x= \int 0\\


xy - \frac{y^2}{2} + x^2 - c_1x = c_2\\


2x^2 - y^2 + 2xy - 2c_1= 2c_2\\

z = 0, y = 2x\\

x = t, y = 2t, z = 0\\

t + 2t + 0 = c_1, c_1 = 3t\\


2t^2 - 4t^2 + 4t^2 - 2(3t) = 2c_2\\

2t^2 - 6t = 2c_2 \\

t^2 - 3t = c_2\\


c_2 = t^2 - c_1\\


2x^2 - y^2 + 2xy - 2c_1= 2(t^2 - c_1)\\

2x^2 - y^2 + 2xy = 2t^2\\


2x^2 - y^2 + 2xy = xy\\

2x^2 - y^2 + xy = 0\\"